What electric field strength would store 11.5 J of energy in every 8.00 mm3 of space?

Two stationary positive point charges, charge 1 of magnitude 3.65 nC and charge 2 of magnitude 1.60 nC , are separated by a distance of 60.0 cm . An electron is released from rest at the point midway between the two charges, and it moves along the line connecting the two charges. What is the speed vfinal of the electron when it is 10.0 cm from charge 1?
Answer #1

1] Let the field be E,

Energy = energy density *Volume

11.5 = 0.5*e0*E^2 *8e-9

E = 1.8*10^10 N/C answer

2] Kinetic energy KE = Decrease in PE = q*decrease in potential

0.5mv^2    = q*k*[q1/r1+q2/r1 - q1/r2 - q2/r2]

0.5*9.1e-31*v^2 = -1.6e-19*9e9*[3.65e-9/0.30 +1.6e-9/0.30 - 3.65e-9/0.10 +1.6e-9/0.50]

v = 7.07*10^6 m/s answer